Problem Solving: Given the following information, what is the WACC?

Commom Stock: 1 million shares outstanding, $40 per share, $1 par value, beta = 1.3; 10,000 bonds outstanding, $1,000 face value each, 8% annual coupon, 22 years to maturity, market price = $1,101.23 per bond

Market risk Premium =         8.6 %, risk-free rate = 5%, marginal tax rate = 35%
Answer #1

Cost of equity:

Information provided:

Risk free rate= 5%

Market risk premium= 8.6%

Beta= 1.3

The cost of equity is calculated using the Capital Asset Pricing Model (CAPM) which is calculated using the formula below:

Ke=Rf+b[E(Rm)-Rf]

Where:

Rf=risk-free rate of return which is the yield on default free debt like treasury notes

Rm=expected rate of return on the market.

b= stock’s beta

Ke= 5% + 1.3*(8.6% - 5%)

     = 5% + 4.68

     = 9.68%.

Cost of debt:

Information provided:

Face value= future value= $1,000

Coupon rate= 8%

Coupon payment= 0.08*1,000= 480

Time= 22 years

Market price= present value= $1101.23

The cost of debt is calculated by computing the yield to maturity.

The yield to maturity is calculated by entering the below in a financial calculator:

FV= 1,000

PV= -1101.23

N= 22

PMT= 80

Press the CPT key and I/Y to compute the yield to maturity.

The value obtained is 7.0788.

Therefore, the before tax cost of debt is 7.08%.

Debt in the capital structure= 10,000*$1101.23= $11,012,300

Equity in the capital structure= 1,000,000*440= $40,000,000

Total firm capital= $11,012,300 + $40,000,000= $51,012,300.

Weight of debt in the firm’s capital = $11,012,300/ $51,012,300

                                                                    = 0.2159*100

                                                                    = 21.59%

Weight of equity in the firm’s capital= $40,000,000/ $51,012,300

                                                                      = 0.7841*100

                                                                      = 78.41%

The weighted average cost of capital is calculated using the below formula:

WACC=Wd*Kd(1-t) +We*Ke

where:

Wd= Percentage of debt in the capital structure.

Kd= The before tax cost of debt

We=Percentage of common stock in the capital structure

Ke= The cost of common stock

T= Tax rate

WACC= 0.2159*7.08%*(1 – 0.35) + 0.7841*9.68%

            = 0.9936% + 7.5901%

            = 8.5837%\rightarrow     8.58%.

In case of any query, kindly comment on the solution